Anna bought two melons and one pineapple for $8.96. If she had purchased one melon and two pineapples, she would have spent $1.49 more. What is the price of a melon? of a pineapple?
Answer by htmentor(1343)

Anna bought two melons and one pineapple for $8.96. If she had purchased one melon and two pineapples, she would have spent $1.49 more. What is the price of a melon? of a pineapple?
========================================================
Let m = the cost of a melon
Let p = the cost of a pineapple
Then we can write the following two equations:
2m + p = 8.96 [1]
m + 2p = 8.96 + 1.49 = 10.45 [2]
Solve for p in [1] and substitute into [2]:
p = 8.96 - 2m
m + 2(8.96-2m) = 10.45
Solve for m:
m + 17.92 - 4m = 10.45
3m = 7.47
m = 2.49
Then p = 8.96 - 2(2.49) = 3.98
So the cost of a melon is $2.49 and the cost of a pineapple is $3.98